# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@ 4 Estimating the Gravitational Wave Background with Horizon-AGN
In this section, we present the methodology for estimating the GW strain spectrum from the population of MBHBs extracted from the simulation Horizon-AGN. We computed it by assuming both circular and eccentric ensembles of MBHBs and investigate the properties of the resulting backgrounds.
@@ 4.1.1 The analytic average
The gravitational wave background (GWB) can be described at each observed frequency $f$ by the characteristic strain $h_{\rm c}(f)$, which can be evaluated by integrating the comoving density of coalescing binaries (Phinney, 2001):
@@
A MBHB on (almost) circular orbit emits GWs at twice its orbital frequency, as higher order modes can be safely ignored since they are suppressed by a factor of $v_{\rm{orb}}/c\ll 1$, where $v_{\rm{orb}}$ is the orbital velocity. In the single dominant mode case we have a one-to-one relationship between the GW frequency in the source frame $f_{\rm s}$ and the binary orbital frequency $f_{\mathrm{orb}}$: ${\rm d}E_{\mathrm{GW}}/{\rm d}\ln f_{\rm s}={\rm d}E_{\mathrm{GW}}/{\rm d}\ln f_{\mathrm{orb}}$. However, this is no longer true for binaries in eccentric orbits. In this case, the emitted GW power is distributed across a set of orbital frequency harmonics $m$, and the distribution depends on the orbital eccentricity of the binary. Loss of orbital momentum through gravitational radiation leads to orbital circularization. As a result, the total GW energy $\text{d}E_{\mathrm{GW}}$ emitted within an observer frequency band centered on $f_{\rm s}$ will include contributions from different harmonics $m$ reflecting different stages of the binary evolution, where $f_{\mathrm{orb}}=f_{\rm s}/m$. In most general (eccentric orbits) case the energy release can be written as (Enoki & Nagashima, 2007):
@@
Here, we introduced the time to coalescence $\tau_{\rm c}$ (measured in the binary rest frame), and the GW luminosity [^2] The influence of MBH spins on the GW luminosity can be safely neglected for such broad orbits and over the observational period of several decades (Mingarelli et al., 2012). of a binary in a circular orbit corresponding to the mean eccentric motion:
@@
where $\mathcal{M}_{\rm c}=(M_{1}M_{2})^{3/5}/(M_{12})^{1/5}$ is the binary chirp mass. The GW spectrum is then obtained by including the $g(m,e)$ function, which gives the relative average (over one orbit) GW power emitted at a given harmonic $m$ by a binary with orbital eccentricity $e$ (see Equation A1 in Peters & Mathews, 1963 for an explicit expression of $g(m,e)$ in terms of Bessel functions).
The energy emitted per logarithmic frequency band is given as a product of GW luminosity and the time the binary spent in that band $\frac{{\rm d}\tau_{\rm c}}{{\rm d}\ln f_{\mathrm{orb}}}$. As mentioned, we use the energy release averaged over the orbit; however, for eccentric motion, it varies significantly over the orbit, where most of the GW power is emitted during the periapse passage. The uneven energy release over time in highly eccentric binaries with long orbital periods could induce non-stationarity in the GWB signal (Falxa et al., 2024). In this work we neglect this effect and consider the averaged (over period) GWB spectrum.
For circular orbits, as described in Section 4.1.3 below, we use the $e\to 0$ limit of Equation 21 directly. For eccentric binaries, we approximate this expression by computing the MBHB dynamics over a finite-width orbital log-frequency grid using the method presented in Section 4.1.4:
@@
that is the set of orbital frequency harmonics, $m$, that would lead to an observed GW frequency within the log-frequency band $\ln f-\Delta\ln f/2,\ln f+\Delta\ln f/2]$ for a binary at redshift $z$.
@@
The sum is over the total energy emitted in GWs by the simulation mergers labeled by $j$, $E_{\mathrm{GW}}^{(j)}$, which can be computed either directly using Equation 21, if an analytic expression for ${\rm d}\tau_{\rm c}/{\rm d}\ln f_{\mathrm{orb}}$ is available, or approximately by Equation 23. Let us emphasize that this expression is fully deterministic, since we compute only one dynamical evolution for each delayed merger in Horizon-AGN.
@@ 4.1.2 Building realizations of the Universe
The analytic computation of the characteristic strain spectrum presented in the previous section is, by definition, deterministic and does not introduce any cosmic variance inherent to the discreteness of the binary population (see Sesana et al., 2008). To estimate the variation in the strain signal due to different realizations [^3] For a given universe, this variance also arises depending on the observer’s location. of the Universe, one must consider the population of inspiralling MBHBs instead of mergers, which are the actual GW sources in the PTA band. In this case, the GWB characteristic strain can be computed as the integral over the redshift of the distribution of inspiralling MBHBs:
@@ 4.1.3 The circular ensemble modeling
In this section, we derive the GWB spectrum for the ensemble of circular MBHBs with the orbital dynamics driven by GW emission and loss-cone scattering. With this simplification, we can carry out all the derivations analytically and use them as a reference for comparison with a more realistic model that allows for eccentric MBHBs.
For circular orbits ($e=0$), the GW power is emitted at the harmonic $m=2$, for which $g(2,0)=1$ and $f_{\rm s}=2f_{\mathrm{orb}}=(1+z)f$. This significantly simplifies Equation 21, as the sum over $m$ is now reduced to a single term ($m=2$). The remaining term is inherently dependent on the binary orbital dynamics, which can be characterized by the evolution of the semi-major axis $a$:
@@
In addition to the GW emission, we incorporate the orbit hardening due to loss-cone scattering (see Equation 7). In the binary rest frame, the dynamics can be modeled as
@@
Here, we neglect the impact of loss-cone scattering on the binary eccentricity evolution, as we force the binaries to remain circular [^4] As shown in figure 1 of Kelley et al. (2017b), loss-cone scattering does not rise eccentricity above 0.2 for most MBHBs in the PTA band, if the binaries were initially circular.. We compute the second term of Equation 31 by assuming that the binaries follow Keplerian orbits.
Figure 1: Examples of the residence time (bottom panel), $\Delta\tau_{\rm c}$, per $\log_{10}$ orbital frequency bin are shown for three binaries from the Horizon-AGN catalog, with chirp masses of $10^{7}$, $10^{8}$, and $10^{9}\,{\rm M_{\odot}}$. We also show their respective eccentricity evolution in the top panel. Eccentric systems are more efficient in dissipating the energy through GWs and this can be seen in the residence time (low panel) as compare to (nearly) circular binaries. The evolution of binaries at high frequencies (starting from a few nHz) is determined by the gravitational radiation which is more efficient in energy dissipation for heavy and eccentric binaries. We also observe a fast circularization of eccentric systems in the top panel.
As a result, Equation 31 can be integrated on a given finite-width orbital frequency grid and plugged in Equation 29 to build universe realizations. In the following, we used log-orbital frequency grids corresponding, for each merger, to 125 log-observer frequency bins equally spaced between $f_{\rm min}=0.8$ nHz and $f_{\rm max}=60$ nHz. The upper bound is constrained by neglecting the frequency evolution (over PTA observing duration) of binaries in our study.
@@ 4.1.4 The eccentric ensemble modeling
As mentioned in section 3, a more detailed description of the MBHB dynamics requires the inclusion of additional physics (viscous drag, eccentricity evolution, etc.) which makes technically not feasible the derivation of analytical expressions for the orbital parameters ($f_{\mathrm{orb}},e,$...) as a function of time. To overcome this limitation, as detailed in section 3, we numerically integrate the dynamics of each merger $j$ and store their orbital parameters on a logarithmic (observer) orbital frequency grid $\log_{10}f_{\mathrm{orb}}^{(k)}=\log_{10}(0.01\,\mathrm{nHz})+k\log_{10}(50\,\mathrm{nHz}/0.01\,\mathrm{nHz})$, for $k\in[0,40]$. One must go to such low orbital frequency to have precise estimate of the GW signal in the PTA band because the GWs emitted by eccentric binaries could have a significant power at frequencies as high as $100\;f_{\mathrm{orb}}$ for $e\gtrsim 0.9$ (Peters & Mathews, 1963). We can extract the residence time in each orbital frequency bin $k$ from the numerical integration and evaluate Equation 29 for each merger $j$. An example for three Horizon-AGN binaries is shown in Figure 1. The eccentricity evolution function $e(f_{\mathrm{orb}}^{(k)})$ for each merger can then be used to compute both Equation 23 and Equation 30 to obtain the analytic average GW spectrum using Equation 26 and the universe realization spectrum using Equation 27.
@@ 4.2.1 Properties of the GW background
The characteristic strain spectra for the circular and eccentric ensembles are shown in Figure 2. As expected, assuming circular binaries driven by GW emission, we find a mean characteristic strain spectrum that follows the power-law behavior of Equation 41, with an amplitude at $f_{\mathrm{yr}}$ of $A_{\mathrm{yr}}=3.4\times 10^{-15}$ and a spectral index of $\alpha=-2/3$ as derived in Phinney (2001). The background amplitude for Horizon-AGN is higher than for many other theoretical models. This can be seen in Appendix A of Agazie et al. (2023c) for pre-2023 models, as well as in new models developed by the EPTA collaboration to explain the detected amplitude (EPTA Collaboration et al., 2024a). In fact, Izquierdo-Villalba et al. (2022) already noted that models that reproduce the background amplitude exceed electromagnetic constraints, such as the bright end of the low-redshift AGN luminosity function. In fact, Horizon-AGN overpredicts the AGN luminosity function (Volonteri et al., 2016), explaining why the background amplitude of the simulation is closer to the value measured by PTAs compared to other models. We stress that in this paper we are not trying to “fit” the amplitude of the background, but we are trying to understand what factors influence the background, and explore the properties of the sources within a unified framework.
@@
Since we include interactions with unbound stars when modeling the binary dynamics for the circular population, we do observe a slight bending of the GW spectrum at low frequencies due to the reduction of the residence time at those frequencies (Ravi et al., 2014). The bend in the Horizon-AGN population starts to significantly affect the spectrum at observer frequencies lower than 2 nHz, which corresponds to an observation duration of around $15.8$ years, in good agreement with previous studies (Sesana, 2013; Kelley et al., 2017b). The current PTA experiments might already be sensitive to this effect (Chen et al., 2024).
@@
Figure 2: 68% (dark) and 90% (light) confidence regions for the characteristic strain $h_{\rm c}$ are shown for 2 000 universe realizations drawn from the Horizon-AGN MBHB population, assuming either circular orbits (blue) or non-circular orbits (orange). We also plot their respective deterministic mean expectation value in dotted and dashed line. For the circular population, we overlay as a dashed-dotted line the average spectrum, assuming the binary dynamics are driven by GW emission only. The inclusion of eccentricity introduces a significant signal reduction at frequency lower than $10\,\rm nHz$, due to the reduction of the MBHBs residence time at low orbital frequency.
@@
As already demonstrated in Sesana et al. (2008), and recently investigated on PTA data in Agazie et al. (2025), the realistic universe realization spectra from the circular population could significantly deviate from the power-law behavior with the mean spectral index corresponding to $\alpha=-2/3$. This is related to the stochasticity of the GW signal from the MBHB population: it is stochastic in the limit of a large number of sources emitting in the same frequency bin during the observation period. The number of binaries per frequency bin decreases with increase in the frequency due to the reduction in their residence time (see Figure 1 and Equation 29), leading to complete loss of stochasticity above 100-$200\,\rm nHz$. In the intermediate frequency range of 10-$50\,\rm nHz$, the decrease in the number of contributing sources increases both the variance and skewness [^5] The skewness (third moment) of the strain distribution at a given observer frequency, across universe realizations, quantifies the asymmetry of the distribution about its mean. of the total characteristic strain distribution across universe realizations. This is a well-known effect, already discussed in Jaffe & Backer (2003); Sesana et al. (2008), and later elaborated in Lamb & Taylor (2024). In Figure 2, the skewness of the strain distribution above $10\,\rm nHz$ is evident for the circular population, where approximately 84% of the GWB spectra have amplitudes below the square root of the average squared strain value. This renders the strain distribution at these frequencies strongly non-Gaussian. As shown in the following, this ultimately results in GWB power spectra with steeper power-law slopes compared to the average spectrum.
@@
When eccentricity is included, we observe two main effects on the GWB spectra. First, we observe the expected bending of the GW spectrum for observer frequencies $\lesssim 10\,\rm nHz$, significantly decreasing the background amplitude at a few nHz. This occurs because eccentricity spreads the GW power across multiple orbital frequency harmonics, combined with the fact that the MBHBs gradually circularize due to the emission of GW (Enoki & Nagashima, 2007). The second effect of eccentricity is a slight reduction in the variance and skewness of the GW strain amplitude around $f_{\mathrm{yr}}$ compared to the circular population. This is caused by the input of highly eccentric binaries with a low orbital frequency, which effectively increases the number of contributions and maintains the stochasticity at higher frequencies.
@@
Figure 3: Median and 68% confidence interval of the differential contribution to the total characteristic strain signal, based on 2 000 universe realizations, shown across MBHB redshift bins for both circular and eccentric populations. We compare the results at two observer frequencies: 3.1 nHz (left panel) and 27.6 nHz (right panel). The vertical lines represent the median redshift values at which the cumulative contribution (integrating from low and high redshifts, respectively) reach 5% of the total strain signal for the circular (solid line) and eccentric (dashed line) populations.

### v2

@@ 4 Estimating the gravitational wave background with Horizon-AGN
In this section, we present the methodology for estimating the GW strain spectrum from the population of MBHBs extracted from the simulation Horizon-AGN. We computed it by assuming both circular and eccentric ensembles of MBHBs and investigated the properties of the resulting backgrounds.
@@ 4.1.1 The analytic average
The gravitational wave background (GWB) can be described at each observed frequency, $f$, by the characteristic strain, $h_{\rm c}(f)$, which can be evaluated by integrating the comoving density of coalescing binaries (Phinney, 2001), expressed as
@@
An MBHB on an (almost) circular orbit emits GWs at twice its orbital frequency, as higher order modes can be safely ignored since they are suppressed by a factor of $v_{\rm{orb}}/c\ll 1$, where $v_{\rm{orb}}$ is the orbital velocity. In the single dominant mode case, we have a one-to-one relationship between the GW frequency in the source frame $f_{\rm s}$ and the binary orbital frequency $f_{\mathrm{orb}}$: ${\rm d}E_{\mathrm{GW}}/{\rm d}\ln f_{\rm s}={\rm d}E_{\mathrm{GW}}/{\rm d}\ln f_{\mathrm{orb}}$. However, this is no longer true for binaries in eccentric orbits. In this case, the emitted GW power is distributed across a set of orbital frequency harmonics, $m$, and the distribution depends on the orbital eccentricity of the binary. Loss of orbital momentum through gravitational radiation leads to orbital circularization. As a result, the total GW energy $\text{d}E_{\mathrm{GW}}$ emitted within an observer frequency band centered on $f_{\rm s}$ will include contributions from different harmonics $m$ reflecting different stages of the binary evolution, where $f_{\mathrm{orb}}=f_{\rm s}/m$. In most general (eccentric orbits) case the energy release can be written as (Enoki and Nagashima, 2007), expressed as
@@
Here, we introduce the time to coalescence, $\tau_{\rm c}$, measured in the binary rest frame, and the GW luminosity [^2] The influence of MBH spins on the GW luminosity can be safely neglected for such broad orbits and over the observational period of several decades (Mingarelli et al., 2012). of a binary in a circular orbit corresponding to the mean eccentric motion:
@@
where $\mathcal{M}_{\rm c}=(M_{1}M_{2})^{3/5}/(M_{12})^{1/5}$ is the binary chirp mass. The GW spectrum is then obtained by including the $g(m,e)$ function, which gives the relative average (over one orbit) GW power emitted at a given harmonic $m$ by a binary with orbital eccentricity $e$ (see Eq. A1 in Peters and Mathews, 1963 for an explicit expression of $g(m,e)$ in terms of Bessel functions).
The energy emitted per logarithmic frequency band is given as a product of GW luminosity and the time the binary spent in that band, expressed as $\frac{{\rm d}\tau_{\rm c}}{{\rm d}\ln f_{\mathrm{orb}}}$. As mentioned, we use the energy release averaged over the orbit; however, for eccentric motion, it varies significantly over the orbit, where most of the GW power is emitted during the periapse passage. The uneven energy release over time in highly eccentric binaries with long orbital periods could induce nonstationarity in the GWB signal (Falxa et al., 2025). In this work we neglect this effect and consider the averaged (over period) GWB spectrum.
For circular orbits, as described in Sect. 4.1.3 below, we use the $e\to 0$ limit of Eq. 21 directly. For eccentric binaries, we approximate this expression by computing the MBHB dynamics over a finite-width orbital log-frequency grid using the method presented in Sect. 4.1.4, namely,
@@
which is the set of orbital frequency harmonics, $m$, that would lead to an observed GW frequency within the log-frequency band $\ln f-\Delta\ln f/2,\ln f+\Delta\ln f/2]$ for a binary at a redshift, $z$.
@@
The sum is over the total energy emitted in GWs by the simulation mergers labeled by $j$, $E_{\mathrm{GW}}^{(j)}$, which can be computed either directly using Eq. 21, if an analytic expression for ${\rm d}\tau_{\rm c}/{\rm d}\ln f_{\mathrm{orb}}$ is available, or approximately by Eq. 23. Let us emphasize that this expression is fully deterministic, since we compute only one dynamical evolution for each delayed merger in Horizon-AGN.
@@ 4.1.2 Building realizations of the Universe
The analytic computation of the characteristic strain spectrum presented in the previous section is, by definition, deterministic and does not introduce any cosmic variance inherent to the discreteness of the binary population (see Sesana et al., 2008). To estimate the variation in the strain signal due to different realizations [^3] For a given Universe, this variance also arises depending on the observer’s location. of the Universe, one must consider the population of inspiralling MBHBs instead of mergers, which are the actual GW sources in the PTA band. In this case, the GWB characteristic strain can be computed as the integral over the redshift of the distribution of inspiralling MBHBs,
@@ 4.1.3 The circular ensemble modeling
In this section, we describe our derivation of the GWB spectrum for the ensemble of circular MBHBs with the orbital dynamics driven by GW emission and loss-cone scattering. With this simplification, we can carry out all the derivations analytically and use them as a reference for comparison with a more realistic model that allows for eccentric MBHBs.
For circular orbits ($e=0$), the GW power is emitted at the harmonic $m=2$, for which $g(2,0)=1$ and $f_{\rm s}=2f_{\mathrm{orb}}=(1+z)f$. This significantly simplifies Eq. 21, as the sum over $m$ is now reduced to a single term ($m=2$). The remaining term is inherently dependent on the binary orbital dynamics, which can be characterized by the evolution of the semi-major axis $a$:
@@
In addition to the GW emission, we incorporate the orbit hardening due to loss-cone scattering (see Eq. 7). In the binary rest frame, the dynamics can be modeled as
@@
Here, we neglected the impact of loss-cone scattering on the binary eccentricity evolution, as we forced the binaries to remain circular [^4] As shown in figure 1 of Kelley et al. (2017a), loss-cone scattering does not raise eccentricity above 0.2 for most MBHBs in the PTA band, if the binaries were initially circular.. We can compute the second term of Eq. 31 by assuming that the binaries follow Keplerian orbits.
Figure 1: Examples of the residence time (bottom panel), $\Delta\tau_{\rm c}$, per $\log_{10}$ orbital frequency bin are shown for three binaries from the Horizon-AGN catalog, with chirp masses of $10^{7}$, $10^{8}$, and $10^{9}\,{\rm M_{\odot}}$. We also show their respective eccentricity evolution in the top panel. Eccentric systems are more efficient in dissipating the energy through GWs and this can be seen in the residence time (low panel) as compared to (nearly) circular binaries. The evolution of binaries at high frequencies (starting from a few nHz) is determined by the gravitational radiation, which is more efficient in energy dissipation for heavy and eccentric binaries. We also observe a fast circularization of eccentric systems in the top panel.
As a result, Eq. 31 can be integrated on a given finite-width orbital frequency grid and plugged in Eq. 29 to build Universe realizations. In the following, we used log-orbital frequency grids corresponding, for each merger, to 125 log-observer frequency bins equally spaced between $f_{\rm min}=0.8$ nHz and $f_{\rm max}=60$ nHz. The upper bound is constrained by neglecting the frequency evolution (over PTA observing duration) of binaries in our study.
@@ 4.1.4 The eccentric ensemble modeling
As mentioned in Sect. 3, a more detailed description of the MBHB dynamics requires the inclusion of additional physics (viscous drag, eccentricity evolution, etc.) which makes it technically unfeasible to obtain a derivation of the analytical expressions for the orbital parameters ($f_{\mathrm{orb}},e,$...) as a function of time. To overcome this limitation (as detailed in Sect. 3), we numerically integrated the dynamics of each merger, $j$, and stored their orbital parameters on a logarithmic (observer) orbital frequency grid $\log_{10}f_{\mathrm{orb}}^{(k)}=\log_{10}(0.01\,\mathrm{nHz})+k\log_{10}(50\,\mathrm{nHz}/0.01\,\mathrm{nHz})$, for $k\in[0,40]$. One must go to such a low orbital frequency to have a precise estimate of the GW signal in the PTA band because the GWs emitted by eccentric binaries could have a significant power at frequencies as high as $100\;f_{\mathrm{orb}}$ for $e\gtrsim 0.9$ (Peters and Mathews, 1963). We can extract the residence time in each orbital frequency bin $k$ from the numerical integration and evaluate Eq. 29 for each merger, $j$. An example for three Horizon-AGN binaries is shown in Fig. 1. The eccentricity evolution function $e(f_{\mathrm{orb}}^{(k)})$ for each merger can then be used to compute both Eq. 23 and Eq. 30 to obtain the analytic average GW spectrum using Eq. 26 and the Universe realization spectrum using Eq. 27.
@@ 4.2.1 Properties of the GW background
The characteristic strain spectra for the circular and eccentric ensembles are shown in Fig. 2. As expected, assuming circular binaries driven by GW emission, we find a mean characteristic strain spectrum that follows the power-law behavior of Eq. 41, with an amplitude at $f_{\mathrm{yr}}$ of $A_{\mathrm{yr}}=3.4\times 10^{-15}$ and a spectral index of $\alpha=-2/3$ as derived in Phinney (2001). The background amplitude for Horizon-AGN is higher than for many other theoretical models. This can be seen in Appendix A of Agazie et al. (2023c) for pre-2023 models, as well as in new models developed by the EPTA collaboration to explain the detected amplitude (EPTA Collaboration et al., 2024a). In fact, Izquierdo-Villalba et al. (2022) already noted that models that reproduce the background amplitude exceed electromagnetic constraints, such as the bright end of the low-redshift AGN luminosity function. In fact, Horizon-AGN overpredicts the AGN luminosity function (Volonteri et al., 2016), explaining why the background amplitude of the simulation is closer to the value measured by PTAs compared to other models. We stress that in this paper we are not trying to “fit” the amplitude of the background, but we are trying to understand what factors influence the background, and explore the properties of the sources within a unified framework.
@@
Since we included interactions with unbound stars when modeling the binary dynamics for the circular population, we did observe a slight bending of the GW spectrum at low frequencies due to the reduction of the residence time at those frequencies (Ravi et al., 2014). Overall, the bend in the Horizon-AGN population starts to significantly affect the spectrum at observer frequencies lower than 2 nHz, which corresponds to an observation duration of around $15.8$ years, in good agreement with previous studies (Sesana, 2013; Kelley et al., 2017a). The current PTA experiments might already be sensitive to this effect (Chen et al., 2024).
@@
Figure 2: 68% (dark) and 90% (light) confidence regions for the characteristic strain $h_{\rm c}$ are shown for 2 000 Universe realizations drawn from the Horizon-AGN MBHB population, assuming either circular orbits (blue) or noncircular orbits (orange). We also plot their respective deterministic mean expectation values in dotted and dashed lines. For the circular population, we overlay as a dashed-dotted line the average spectrum, assuming the binary dynamics are driven by GW emission only. The inclusion of eccentricity introduces a significant signal reduction at frequencies lower than $10\,\rm nHz$, due to the reduction of the MBHBs residence time at low orbital frequency.
@@
As already demonstrated in Sesana et al. (2008) and recently investigated on the basis of PTA data in Agazie et al. (2025), the realistic Universe realization spectra from the circular population could significantly deviate from the power-law behavior with the mean spectral index corresponding to $\alpha=-2/3$. This is related to the stochasticity of the GW signal from the MBHB population: it is stochastic in the limit of a large number of sources emitting in the same frequency bin during the observation period. The number of binaries per frequency bin decreases with an increase in the frequency due to the reduction in their residence time (see Fig. 1 and Eq. 29), leading to a complete loss of stochasticity above 100-$200\,\rm nHz$. In the intermediate frequency range of 10-$50\,\rm nHz$, the decrease in the number of contributing sources increases both the variance and skewness [^5] The skewness (third moment) of the strain distribution at a given observer frequency, across Universe realizations, quantifies the asymmetry of the distribution about its mean. of the total characteristic strain distribution across Universe realizations. This is a well-known effect, already discussed in Jaffe and Backer (2003); Sesana et al. (2008), and later elaborated in Lamb and Taylor (2024). In Fig. 2, the skewness of the strain distribution above $10\,\rm nHz$ is evident for the circular population, where approximately 84% of the GWB spectra have amplitudes below the square root of the average squared strain value. This renders the strain distribution at these frequencies strongly non-Gaussian. As shown in the following, this ultimately results in GWB power spectra with steeper power-law slopes compared to the average spectrum.
@@
When eccentricity is included, we observe two main effects on the GWB spectra. First, we observe the expected bending of the GW spectrum for observer frequencies $\lesssim 10\,\rm nHz$, significantly decreasing the background amplitude at a few nanohertz (nHz). This occurs because eccentricity spreads the GW power across multiple orbital frequency harmonics, combined with the fact that the MBHBs gradually circularize due to the emission of GW (Enoki and Nagashima, 2007). The second effect of eccentricity is a slight reduction in the variance and skewness of the GW strain amplitude around $f_{\mathrm{yr}}$ compared to the circular population. This is caused by the input of highly eccentric binaries with a low orbital frequency, which effectively increases the number of contributions and maintains the stochasticity at higher frequencies.
@@
Figure 3: Median and 68% confidence interval of the differential contribution to the total characteristic strain signal, based on 2 000 Universe realizations, shown across MBHB redshift bins for both circular and eccentric populations. We compare the results at two observer frequencies: 3.1 nHz (left panel) and 27.6 nHz (right panel). The vertical lines represent the median redshift values at which the cumulative contribution (integrating from low and high redshifts, respectively) reaches 5% of the total strain signal for the circular (solid line) and eccentric (dashed line) populations.